### v1

# Polarization-selective quantum cooperative response in dual-species atom arrays

Abstract
Atom arrays have emerged as a powerful platform for quantum light-matter interfaces, yet single-species arrays are constrained by in-plane symmetry, restricting polarization control. Here we investigate the cooperative optical response of dual-species subwavelength atom arrays, in which intrinsic polarizability difference breaks in-plane symmetry. By engineering the lattice spacing and detunings, the arrays exhibit polarization-dependent subradiant modes, enabling complete reflection of specific polarization component. Leveraging this mechanism, we assemble array units as functional pixels and demonstrate a scalable polarization-selective quantum light modulator. Our work establishes a dynamically reconfigurable atomic-photonic platform for versatile subwavelength quantum optical elements.
journal: opticajournal
1 Introduction
Atom arrays play a central role in quantum science and technology, offering a highly controllable and versatile platform for frontier areas including quantum computation, quantum simulation, quantum metrology, and quantum communication [1, 2, 3, 4]. In particular, ordered atom arrays with subwavelength spacing provide unique advantages in quantum optics and quantum information processing by enabling strong light-atom interactions and cooperative optical responses [5]. Recent studies have been focusing on cooperative quantum phenomena such as superradiance and subradiance [6, 7, 8, 9], highlighting the potential of atom arrays for quantum information processing, photon storage, and engineered quantum optical elements [8, 9, 10, 11, 5].
Theoretically, two-dimensional subwavelength atom arrays support cooperative scattering resonances at specific magic lattice constants, enabling near-unity extinction of incident light [12, 13, 14, 15]. By tailoring array geometries and multilevel atomic configurations, a broad range of optical functionalities have been proposed, including Huygens surfaces, negative refraction, and linear optical elements [16, 17, 18, 19]. Experimentally, two-dimensional arrays containing a few hundred atoms have been realized as highly efficient optical mirrors [20], and further extended to quantum metasurfaces with dynamic control enabled by auxiliary Rydberg atoms and electromagnetically induced transparency [10, 21]. While these single-species atom arrays offer foundational insights, their inherent symmetry restricts selective control over polarization, suggesting a need for more complex, multi-species configurations.
Recently, dual-species atom arrays have shown distinct advantages in quantum computing, where different atomic species can act as data and auxiliary qubits, enabling parallel quantum operations and fully connected networks [22, 23, 24]. More importantly, intrinsic differences in atomic polarizability introduce additional tunable degrees of freedom, opening new possibilities for multi-degree-of-freedom optical field manipulation.
In this work, we systematically study the cooperative optical response of dual-species subwavelength atom arrays and identify polarization-dependent subradiant modes accompanied by pronounced optical anisotropy. We reveal the physical origin of polarization selectivity near a symmetric detuning point, which arises from strong coherent scattering with balanced radiative losses. Leveraging this mechanism, we design and demonstrate a scalable polarization-selective quantum light modulator composed of multiple dual-species array pixels, and discuss feasible experimental implementations. By bridging fundamental physical insights and device-level architectures, our work opens new avenues for realizing multifunctional quantum light–matter interfaces.
2 Dual-species atom array
We consider the cooperative optical response of a two-dimensional subwavelength square array of lattice constant $a$ containing two atomic species arranged in a stripe pattern [Fig. 1(a)]. Each atom is modeled as a two-level system with transition frequency $\omega_{A}$ or $\omega_{B}$. A light beam incident along the $z$-axis illuminates the array in the xxyy plane at $z=0$. Both species are driven by light of frequency $\omega$, with detunings $\delta_{A}=\omega-\omega_{A}$ and $\delta_{B}=\omega-\omega_{B}$, as shown in Fig. 1(b). We choose the two species to be isotopes so that they share similar energy levels and can be driven near resonance by the same field, with $\omega$ close to both $\omega_{A}$ and $\omega_{B}$.
Figure 1: (a) Array of two isotopic atomic species, with blue and red spheres denoting atoms of type A and B, respectively. (b) Corresponding two-level energy structure.
The incident light is scattered by the atom array with each atom acting a quantum emitter, which can be modeled as pointlike particles with linear isotropic polarizabilities [13, 12]. We consider the cooperative response by analyzing the total electric field, which is the sum of the incident and scatterd field and can be obtained by solving the wave equation, $\nabla\times\nabla\times\mathbf{E}-k^{2}\mathbf{E}=k^{2}\mathbf{P}/\varepsilon_{0}$, where $\mathbf{P}$ is the total polarization, $\varepsilon_{0}$ is the vacuum permittivity, $k=\omega/c$ is the wave number and $c$ is the light speed. The general solution can be expressed as
$E_{\nu}(\mathbf{r})=E_{\mathrm{inc},\nu}(\mathbf{r})+\frac{k^{2}}{\varepsilon_{0}}\sum_{\mu}\int\mathrm{d}\mathbf{r}^{\prime}G_{\nu\mu}(k,\mathbf{r},\mathbf{r}^{\prime})P_{\mu}(\mathbf{r}^{\prime}),$ (1)
where $\nu,\mu=x,y,z$ denotes the spatial components. $\mathbf{E}_{\mathrm{inc}}(\mathbf{r})$ represents the incident field. $G_{\nu\mu}(k,\mathbf{r},\mathbf{r}^{\prime})$ is the component of dyadic Green’s function, describing the fields produced by an infinitesimal electric current source (see Supplement 1 for explicit expression).
Considering the linear response regime, $\mathbf{P}$ is proportional to the electric field. For the discrete atom array $\mathbf{P}(\mathbf{r})=\sum_{n}\alpha_{n}\mathbf{E}(\mathbf{r})\delta(\mathbf{r}-\mathbf{r}_{n})$, where $n$, $\mathbf{r}_{n}$, and $\alpha_{n}$ denote the index, position, polarizability, respectively. Substituting into Eq. (1) gives the electric field
$E_{\nu}(\mathbf{r})$ $=E_{\mathrm{inc},\nu}(\mathbf{r})+\frac{4\pi^{2}}{\varepsilon_{0}\lambda^{2}}\underset{\mu}{\sum}\sum_{n}\alpha_{n}G_{\nu\mu}(k,\mathbf{r},\mathbf{r}_{n})E_{\mu}(\mathbf{r}_{n}),$ (2)
where $\lambda=2\pi/k$ is the wavelength. For a closed two-level $J=0$ to $J=1$ atomic transition [25, 26], the polarizability takes the form $\alpha_{n}=-\frac{3\varepsilon_{0}\lambda^{3}}{4\pi^{2}}\frac{\gamma/2}{\delta_{n}+\mathrm{i}\gamma/2}$ for small detuning. For atoms of type A or B, the subscript $n$ is replaced by A or B. The spontaneous emission rate $\gamma$ is taken to be the same for both isotopes (see Supplement 1). Equation (2) shows that the local field is the sum of the incident field and the fields scattered by all other emitters. Evaluated at the atomic positions, it becomes a self-consistent set of equations that can be solved numerically for arbitrary incident fields (see Supplement 1).
To quantify the scattering of the dual-species array, we define the transmission coefficient by computing the total power collected by a downstream lens. The lens is placed at $z=z_{L}=150\lambda$ centered on $x=y=0$, with radius $R_{L}=90\lambda$, large enough to avoid finite-size effects while maintaining a realistic numerical aperture ($\mathrm{NA}=R_{L}/z_{L}=0.6$). The collected power is $P=(\varepsilon_{0}c^{2}/2)\int_{\mathcal{L}}\mathrm{Re}(\mathbf{E}\times\mathbf{B}^{\ast})\cdot\mathrm{d}\mathbf{A}$, where $\mathcal{L}$ is the lens area and $\mathrm{d}\mathbf{A}=\mathrm{d}A\mathbf{e}_{z}$. For a plane wave, the magnetic field is $\mathbf{B}=\mathbf{e}_{z}\times\mathbf{E}/c$. In the absence of atom array, the collected power is $P_{\mathrm{inc}}$. Then, the transmission is defined as $T=P/P_{\mathrm{inc}}$. Because the stripe pattern breaks the symmetry between the $x$ and $y$ directions, the transmission generally differs for horizontal ($x$) and vertical ($y$) polarizations. This polarization dependence provides an additional degree of freedom for manipulating subradiant effects, enhancing the functionality of cooperative atom arrays as quantum light–matter interfaces.
3 Cooperative optical response
Figure 2: Transmissions $T_{x}$ and $T_{y}$ as functions of lattice constant $a$. (a,b) The detuning is fixed at $\delta_{A}/\gamma=1$, while $\delta_{B}=1,2,5,10,100,\infty$. (c,d) The detuning is fixed at $\delta_{B}/\gamma=-1$, while $\delta_{A}=1,2,5,10,100,\infty$. Different curve colors denote different values of $\delta_{A}$ or $|\delta_{B}|$, as indicated in the legend of (a). Calculations assume an $N=26\times 26$ atom array and a diagonally polarized incident Gaussian beam with waist $w_{0}=0.3\sqrt{N}a$.
In dual-species atom arrays, two detunings $\delta_{A}$ and $\delta_{B}$ can be independently tuned to control the atomic polarizabilities and, together with the lattice constant, determine the cooperative optical response. We first consider the case where the incident light is near resonant with one atomic species. Specifically, we fix $\delta_{A}$ and vary $\delta_{B}$, and numerically calculate the transmissions $T_{x}$ and $T_{y}$ for horizontally ($x$) and vertically ($y$) polarized incident light. We define $T_{\mu}=P_{\mu}/P_{\mathrm{inc},\mu}$ with $\mu=x,y$, where $P_{\mu}$ and $P_{\mathrm{inc},\mu}$ are the transmitted and incident optical powers in the $\mu$-polarized component, respectively.
The resulting $T_{x}$ and $T_{y}$ as functions of $a$ are shown in Fig. 2 (a,b). For sufficiently large $|\delta_{B}|$, type B atoms become effectively transparent, and the dual-species square array reduces to an effective single-species rectangular array with lattice constants $a_{x}=a$ and $a_{y}=2a$, corresponding to the curves with $\delta_{B}=\infty$ (see Supplement 1 for detailed analysis). For smaller $|\delta_{B}|$, both $T_{x}$ and $T_{y}$ exhibit transmission zeros at distinct lattice constants, indicating polarization-dependent subradiant modes. At even smaller $|\delta_{B}|$, $T_{y}$ shows two transmission zeros, reflecting enhanced collective response when both atomic species are near resonance. Overall, a pronounced anisotropy between the two orthogonal polarizations is observed.
A slightly different picture occurs when $\delta_{B}$ is fixed and $\delta_{A}$ is varied, as shown in Fig. 2 (c,d). Here, the transition of type B atoms is red-detuned with $\delta_{B}=-1$. For sufficiently large $\delta_{A}$, the array again reduces to an effective single-species rectangular array, corresponding to the curves with $\delta_{A}=\infty$. At smaller $\delta_{A}$, both $T_{x}$ and $T_{y}$ exhibit polarization-dependent subradiant modes, also indicating an enhanced collective response when both atomic species are near resonance.
When both detunings $\delta_{A}$ and $\delta_{B}$ are not too large, the dual-species array exhibits strong cooperative effects, leading to substantial suppression of transmission. We then consider the case where the light frequency lies between the two transition frequencies $\omega_{A}$ and $\omega_{B}$. A particularly symmetric choice is the midpoint, $\delta_{A}=-\delta_{B}=\delta$, for which $\mathrm{Re}(\alpha_{A})=-\mathrm{Re}(\alpha_{B})$ and $\mathrm{Im}(\alpha_{A})=\mathrm{Im}(\alpha_{B})$. In this configuration, the two atom species have balanced radiative loss, creating an exotic interference environment that enables strong collective effects through coherent scattering.
Figure 3: The transmissions and subradiant modes. (a, b) Transmissions $T_{x}$ and $T_{y}$ as functions of the lattice constant $a$. Colors represent $\delta/\gamma=0.1,0.3,0.5,0.7,0.9$ (see legend in (a)). (c, d) Lattice constants yielding zero transmission versus detuning $\delta$, corresponding to $T_{x}$ and $T_{y}$, respectively. Blue solid and red dotted lines denote the first and second zeros (legend in (d)); The inset of (c) extends the results to a wider range of $\delta/\gamma$. The black straight line indicates the asymptotic scaling $a\sim\delta^{-0.32}$ at large $\delta$. Calculations assume the same atom array and incident beam as in Fig. 2.
For different values of $\delta$, we numerically calculate the transmissions $T_{x}$ and $T_{y}$ as functions of lattice constant $a$, extracted from the electric fields solved via the self-consistent equation [Eq. (2)]. In Fig. 3 (a) and (b), we plot $T_{x}$ and $T_{y}$ for various detunings $\delta$. At $\delta=0$, the array effectively reduces to a single-species array, where $T_{x}=T_{y}$ and zero transmission occurs at $a/\lambda\approx 0.2$, or $0.8$, corresponding to cooperative resonances, consistent with previous single-species square array calculations [12, 13].
As $\delta$ increases, the dual-species nature emerges, and $T_{x}$ and $T_{y}$ diverge as the stripe pattern introduces anisotropy between the $x$- and $y$-directions. At larger $\delta$, the number of transmission zeros decreases for both $T_{x}$ and $T_{y}$. The lattice constants for zero transmission as functions of $\delta$ are shown in Fig. 3 (c) and (d). In Fig. 3 (c), for small detunings ($0<\delta<0.11$), $T_{x}$ have two zeros at $a/\lambda\approx 0.2$ and $0.8$. For $\delta>0.11$, one zero around $a/\lambda\approx 0.8$ disappears, leaving only one zero. As $\delta$ increases, the remaining transmission zero shifts to smaller $a$ and approximately follows a power-law scaling at large $\delta$, $a/\lambda\sim 0.22(\delta/\gamma)^{-0.32}$ (see the inset of Fig. 3 (c)). In Fig. 3 (d), similarly at small detunings ($0<\delta<0.07$), $T_{y}$ have two zeros at $a/\lambda\approx 0.2$ and $0.8$. As $\delta$ increases, one zero disappears at $\delta\approx 0.07$, reappears at $\delta\approx 0.26$, and merges with the other zero at $\delta\approx 0.88$. No zero transmission exists for $\delta\gtrsim 1.08$.
The anisotropy between $T_{x}$ and $T_{y}$ provides a novel route to engineer subradiant modes in a polarization-selective manner. By choosing appropriate detuning $\delta$ and lattice constant $a$, one polarization state can enter a subradiant mode with near-complete reflection, while the orthogonal polarization remains transmissive. To further illustrate this effect, Fig. 4 shows the light intensity and polarization distribution for a diagonally polarized incident beam with $\delta/\gamma=0.5$ and $a/\lambda=0.4$. The dual-species array nearly completely reflects the $y$-polarized component with $T_{y}\approx 0$, while producing $x$-polarized light with $T_{x}\approx 0.50$. Remarkably, the transmissive polarization is parallel to the stripe-pattern direction. This demonstrates dual-species atom arrays acting as subwavelength optical polarizers, enabling programmable and independent polarization control at the atomic scale.
Figure 4: (a–c) Total light intensity and polarization distributions at propagation distances $z=14\lambda$, $50\lambda$, and $90\lambda$, respectively. The color bar denotes the normalized optical intensity $I=(|E_{x}|^{2}+|E_{y}|^{2})/|E_{0}|^{2}$, where $E_{0}$ is the peak amplitude of the incident Gaussian beam. White lines indicate the polarization distributions. Calculations assume $a/\lambda=0.4$ and $\delta/\gamma=0.5$. The figures span spatial ranges $x,y\in\left[-10\lambda,10\lambda\right]$.
4 Polarization-selective quantuam light modulator
By introducing an ancillary atom which can be excited to a Rydberg state, the atom array can be switched between fully reflective and transparent regimes via electromagnetically induced transparency and Rydberg blockade [10, 21]. The array thus functions as a quantum metasurface that can exist in a superposition of states reflecting and transmitting the incident photons [10, 21]. In a dual-species atom array, such a quantum metasurface becomes polarization-selective, enabling control over specific polarization states and introducing additional tunable degrees of freedom. Multiple quantum metasurfaces can be further assembled into a quantum light modulator, with each quantum metasurface acting as a pixel, allowing the generation of highly entangled photonic states [10].
Building on this idea, dual-species atom arrays can be used to construct a polarization-selective quantum light modulator, where each pixel selects a desired polarization state. We demonstrate this concept by assembling multiple dual-species atom arrays as pixels, forming a programmable platform for pixel-by-pixel polarization control. Figure 5 (a) shows a superarray composed of $2\times 2$ pixels, where each pixel is a dual-species atom array with a prescribed stripe-pattern direction (see Supplement 1 for additional configurations). The lattice sites between pixels are filled with type-A atoms, serving as isolation regions. At $\delta_{A}/\gamma=0.5$ and lattice constant $a/\lambda=0.4$, a square array of type-A atoms exhibits near-perfect reflection for both polarizations with $T_{x}=T_{y}\approx 0$, thereby suppressing crosstalk between neighboring pixels.
We further consider a superarray arranged similarly to Fig. 5 (a), with a total atom number $N=71\times 71$, where each pixel contains $30\times 30$ atoms and the isolation width is $11$ lattice sites. The spatial distributions of light intensity and polarization are shown in Fig. 5 (b), demonstrating the designed pixel-by-pixel modulation of intensity and polarization. This establishes a foundation for assembling polarization-selective quantum light modulators [10] and highlights dual-species atom arrays as highly tunable quantum photonic systems at the atomic scale, enabling precise manipulation of vectorial optical fields and a quantum light–matter interface beyond the capabilities of conventional subwavelength optical metamaterials [27, 28, 29] .
Figure 5: (a) Patterned superarray with $2\times 2$ pixels, where each pixel is a dual-species stripe-pattern atom array. (b) Total light intensity and polarization distribution at $z=14\lambda$. Calculations assume an $N=71\times 71$ atom array with $a/\lambda=0.4$, and a diagonally polarized incident beam with waist $w_{0}=0.3\sqrt{N}a$ and $\delta/\gamma=0.5$. Each pixel contains $30\times 30$ atoms and the isolation width is $11$ lattice sites. The color bar denotes the normalized optical intensity $I=(|E_{x}|^{2}+|E_{y}|^{2})/|E_{0}|^{2}$. White lines or ellipses indicate the polarization distribution.
5 Experimental Proposal
Atom arrays of ytterbium (Yb) have emerged as a powerful platform for quantum computation and quantum precision measurement. Yb atoms feature a transition from $6s6p\penalty 10000\ {}^{3}P_{0}$ to $5d6s\penalty 10000\ {}^{3}D_{1}$ at a telecom-band wavelength $\lambda\approx 1.4\mu\mathrm{m}$, realizing a $J=0\rightarrow J=1$ two-level system, enabling confinement in optical lattices or tweezer arrays with subwavelength spacing. We consider a dual-species atom array of the isotopes 171Yb and 173Yb. Their transition frequencies are $f_{171}=215869124.9\mathrm{MHz}$ and $f_{173}=215869151.9\mathrm{MHz}$, respectively [30]. Using the spontaneous emission formula $\gamma={d^{2}\omega^{3}}/(3\pi\varepsilon_{0}\hbar c^{3})$, we obtain $\gamma_{171}\approx\gamma_{173}\approx 5.93\mathrm{MHz}$, yielding a separation of $|f_{173}-f_{171}|\sim 4.55\gamma$. This energy offset can be continuously tuned via AC Stark shifts induced by coupling to nearby auxiliary levels, allowing controllable detunings $\delta_{A}$ and $\delta_{B}$ within the parameter range considered in this work.
6 Summary
Our work systematically investigates the cooperative optical response and polarization control mechanisms in dual-species atom arrays. By independently tuning the light detunings and lattice spacing, we realize polarization-dependent subradiant modes and identify strong collective scattering at a symmetric detuning point with balanced radiative losses. Exploiting this mechanism, the array operates as a subwavelength polarizer with atomic-scale polarization selectivity. We further propose and validate a scalable polarization-selective quantum light modulator, enabling spatially programmable polarization control through pixel integration and engineered isolation. These results establish dual-species atom arrays as a highly tunable quantum photonic platform and provide a theoretical foundation for advanced quantum light–matter interfaces.
[…]
Funding S.Z. acknowledges support from the Youth Innovation Promotion Association, Chinese Academy of Sciences (2023399). X.L. acknowledges support from the National Natural Science Foundation of China (U24A6010).
[…]
Disclosures The authors declare no conflicts of interest.
[…]
Data Availability Statement Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.
[…]
Supplemental document See Supplement 1 for supporting content.

### v2

@@
 Here we investigate the cooperative optical response of dual-species subwavelength atom arrays, in which intrinsic polarizability difference breaks in-plane symmetry.
-By engineering the lattice spacing and detunings, the arrays exhibit polarization-dependent subradiant modes, enabling complete reflection of specific polarization component.
+By engineering the lattice constants and detunings, the arrays exhibit polarization-dependent subradiant modes, enabling complete reflection of a specific polarization component.
 Leveraging this mechanism, we assemble array units as functional pixels and demonstrate a scalable polarization-selective quantum light modulator.
@@
 Figure 1: (a) Array of two isotopic atomic species, with blue and red spheres denoting atoms of type A and B, respectively. (b) Corresponding two-level energy structure.
-The incident light is scattered by the atom array with each atom acting a quantum emitter, which can be modeled as pointlike particles with linear isotropic polarizabilities [13, 12].
-We consider the cooperative response by analyzing the total electric field, which is the sum of the incident and scatterd field and can be obtained by solving the wave equation, $\nabla\times\nabla\times\mathbf{E}-k^{2}\mathbf{E}=k^{2}\mathbf{P}/\varepsilon_{0}$, where $\mathbf{P}$ is the total polarization, $\varepsilon_{0}$ is the vacuum permittivity, $k=\omega/c$ is the wave number and $c$ is the light speed.
+The incident light is scattered by the atom array with each atom acting as a quantum emitter, which can be modeled as pointlike particles with linear isotropic polarizabilities [8, 24].
+We consider the cooperative response by analyzing the total electric field, which is the sum of the incident and scattered fields and can be obtained by solving the wave equation, $\nabla\times\nabla\times\mathbf{E}-k^{2}\mathbf{E}=k^{2}\mathbf{P}/\varepsilon_{0}$, where $\mathbf{P}$ is the total polarization, $\varepsilon_{0}$ is the vacuum permittivity, $k=\omega/c$ is the wave number and $c$ is the light speed.
 The general solution can be expressed as
 $E_{\nu}(\mathbf{r})=E_{\mathrm{inc},\nu}(\mathbf{r})+\frac{k^{2}}{\varepsilon_{0}}\sum_{\mu}\int\mathrm{d}\mathbf{r}^{\prime}G_{\nu\mu}(k,\mathbf{r},\mathbf{r}^{\prime})P_{\mu}(\mathbf{r}^{\prime}),$ (1)
-where $\nu,\mu=x,y,z$ denotes the spatial components.
+where $\nu,\mu=x,y,z$ denote the spatial components.
 $\mathbf{E}_{\mathrm{inc}}(\mathbf{r})$ represents the incident field.
@@
 Considering the linear response regime, $\mathbf{P}$ is proportional to the electric field.
-For the discrete atom array $\mathbf{P}(\mathbf{r})=\sum_{n}\alpha_{n}\mathbf{E}(\mathbf{r})\delta(\mathbf{r}-\mathbf{r}_{n})$, where $n$, $\mathbf{r}_{n}$, and $\alpha_{n}$ denote the index, position, polarizability, respectively.
+For the discrete atom array, $\mathbf{P}(\mathbf{r})=\sum_{n}\alpha_{n}\mathbf{E}(\mathbf{r})\delta(\mathbf{r}-\mathbf{r}_{n})$, where $n$, $\mathbf{r}_{n}$, and $\alpha_{n}$ denote the index, position, and polarizability, respectively.
 Substituting into Eq.
@@
 where $\lambda=2\pi/k$ is the wavelength.
-For a closed two-level $J=0$ to $J=1$ atomic transition [25, 26], the polarizability takes the form $\alpha_{n}=-\frac{3\varepsilon_{0}\lambda^{3}}{4\pi^{2}}\frac{\gamma/2}{\delta_{n}+\mathrm{i}\gamma/2}$ for small detuning.
+For a closed two-level $J=0$ to $J=1$ atomic transition [15], the polarizability takes the form $\alpha_{n}=-\frac{3\varepsilon_{0}\lambda^{3}}{4\pi^{2}}\frac{\gamma/2}{\delta_{n}+\mathrm{i}\gamma/2}$ for small detuning.
 For atoms of type A or B, the subscript $n$ is replaced by A or B.
@@
 Evaluated at the atomic positions, it becomes a self-consistent set of equations that can be solved numerically for arbitrary incident fields (see Supplement 1).
-To quantify the scattering of the dual-species array, we define the transmission coefficient by computing the total power collected by a downstream lens.
-The lens is placed at $z=z_{L}=150\lambda$ centered on $x=y=0$, with radius $R_{L}=90\lambda$, large enough to avoid finite-size effects while maintaining a realistic numerical aperture ($\mathrm{NA}=R_{L}/z_{L}=0.6$).
-The collected power is $P=(\varepsilon_{0}c^{2}/2)\int_{\mathcal{L}}\mathrm{Re}(\mathbf{E}\times\mathbf{B}^{\ast})\cdot\mathrm{d}\mathbf{A}$, where $\mathcal{L}$ is the lens area and $\mathrm{d}\mathbf{A}=\mathrm{d}A\mathbf{e}_{z}$.
+To quantify the scattering of the dual-species array, we define the transmission coefficient by calculating the total power collected over a two-dimensional circular area located at $z=z_{L}=150\lambda$, centered at $x=y=0$ in the $x$-$y$ plane, with radius $R_{L}=90\lambda$, which is sufficiently large to avoid finite-size effects [8].
+The collected power is $P=(\varepsilon_{0}c^{2}/2)\int_{\mathcal{L}}\mathrm{Re}(\mathbf{E}\times\mathbf{B}^{\ast})\cdot\mathrm{d}\mathbf{A}$, where $\mathcal{L}$ is the circular area and $\mathrm{d}\mathbf{A}=\mathrm{d}A\mathbf{e}_{z}$.
 For a plane wave, the magnetic field is $\mathbf{B}=\mathbf{e}_{z}\times\mathbf{E}/c$.
@@
 Figure 2: Transmissions $T_{x}$ and $T_{y}$ as functions of lattice constant $a$.
-(a,b) The detuning is fixed at $\delta_{A}/\gamma=1$, while $\delta_{B}=1,2,5,10,100,\infty$.
-(c,d) The detuning is fixed at $\delta_{B}/\gamma=-1$, while $\delta_{A}=1,2,5,10,100,\infty$.
+(a,b) The detuning is fixed at $\delta_{A}/\gamma=1$, while $\delta_{B}/\gamma=1,2,5,10,100,\infty$.
+(c,d) The detuning is fixed at $\delta_{B}/\gamma=-1$, while $\delta_{A}/\gamma=1,2,5,10,100,\infty$.
 Different curve colors denote different values of $\delta_{A}$ or $|\delta_{B}|$, as indicated in the legend of (a).
@@
 2 (a,b).
-For sufficiently large $|\delta_{B}|$, type B atoms become effectively transparent, and the dual-species square array reduces to an effective single-species rectangular array with lattice constants $a_{x}=a$ and $a_{y}=2a$, corresponding to the curves with $\delta_{B}=\infty$ (see Supplement 1 for detailed analysis).
+For sufficiently large $|\delta_{B}|$, type B atoms become effectively transparent, and the dual-species square array reduces to an effective single-species rectangular array with lattice constants $a_{x}=a$ and $a_{y}=2a$, corresponding to the curves with $|\delta_{B}|=\infty$ (see Supplement 1 for detailed analysis).
 For smaller $|\delta_{B}|$, both $T_{x}$ and $T_{y}$ exhibit transmission zeros at distinct lattice constants, indicating polarization-dependent subradiant modes.
@@
 2 (c,d).
-Here, the transition of type B atoms is red-detuned with $\delta_{B}=-1$.
+Here, the transition of type B atoms is red-detuned with $\delta_{B}/\gamma=-1$.
 For sufficiently large $\delta_{A}$, the array again reduces to an effective single-species rectangular array, corresponding to the curves with $\delta_{A}=\infty$.
 At smaller $\delta_{A}$, both $T_{x}$ and $T_{y}$ exhibit polarization-dependent subradiant modes, also indicating an enhanced collective response when both atomic species are near resonance.
-When both detunings $\delta_{A}$ and $\delta_{B}$ are not too large, the dual-species array exhibits strong cooperative effects, leading to substantial suppression of transmission.
+When both detunings $\delta_{A}$ and $\delta_{B}$ are not too large, the dual-species array exhibits strong cooperative effects, leading to substantial suppression of transmission (see Supplement 1 for detailed analysis).
 We then consider the case where the light frequency lies between the two transition frequencies $\omega_{A}$ and $\omega_{B}$.
-A particularly symmetric choice is the midpoint, $\delta_{A}=-\delta_{B}=\delta$, for which $\mathrm{Re}(\alpha_{A})=-\mathrm{Re}(\alpha_{B})$ and $\mathrm{Im}(\alpha_{A})=\mathrm{Im}(\alpha_{B})$.
+A particularly symmetric choice is the midpoint, $\delta_{A}=-\delta_{B}=\delta$.
+With identical spontaneous emission rate $\gamma$, this leads to $\mathrm{Re}(\alpha_{A})=-\mathrm{Re}(\alpha_{B})$ and $\mathrm{Im}(\alpha_{A})=\mathrm{Im}(\alpha_{B})$ (see Supplement 1).
 In this configuration, the two atom species have balanced radiative loss, creating an exotic interference environment that enables strong collective effects through coherent scattering.
@@
 Blue solid and red dotted lines denote the first and second zeros (legend in (d)); The inset of (c) extends the results to a wider range of $\delta/\gamma$.
-The black straight line indicates the asymptotic scaling $a\sim\delta^{-0.32}$ at large $\delta$.
+The black straight line shows a power-law fitting $a\sim\delta^{-0.32}$ at large $\delta$.
 Calculations assume the same atom array and incident beam as in Fig.
@@
 In Fig.
-3 (c), for small detunings ($0<\delta<0.11$), $T_{x}$ have two zeros at $a/\lambda\approx 0.2$ and $0.8$.
-For $\delta>0.11$, one zero around $a/\lambda\approx 0.8$ disappears, leaving only one zero.
-As $\delta$ increases, the remaining transmission zero shifts to smaller $a$ and approximately follows a power-law scaling at large $\delta$, $a/\lambda\sim 0.22(\delta/\gamma)^{-0.32}$ (see the inset of Fig.
-3 (c)).
+3 (c), for small detunings ($0<\delta/\gamma<0.11$), $T_{x}$ has two zeros at $a/\lambda\approx 0.2$ and $0.8$.
+For $\delta/\gamma>0.11$, one zero around $a/\lambda\approx 0.8$ disappears, leaving only one zero.
+As $\delta/\gamma$ increases, the remaining transmission zero shifts toward smaller $a/\lambda$ and, at large $\delta/\gamma$, approximately follows a numerical power-law fitting, $a/\lambda\sim 0.22(\delta/\gamma)^{-0.32}$ (see the inset of Fig.
+3(c)).
 In Fig.
-3 (d), similarly at small detunings ($0<\delta<0.07$), $T_{y}$ have two zeros at $a/\lambda\approx 0.2$ and $0.8$.
-As $\delta$ increases, one zero disappears at $\delta\approx 0.07$, reappears at $\delta\approx 0.26$, and merges with the other zero at $\delta\approx 0.88$.
-No zero transmission exists for $\delta\gtrsim 1.08$.
+3 (d), similarly at small detunings ($0<\delta/\gamma<0.07$), $T_{y}$ has two zeros at $a/\lambda\approx 0.2$ and $0.8$.
+As $\delta$ increases from 0, the system evolves from a single-species array to a dual-species array with increasing anisotropy between the $x$- and $y$-directions (see Supplement 1 for detailed analysis).
+One zero disappears at $\delta/\gamma\approx 0.07$, reappears at $\delta/\gamma\approx 0.26$, and merges with the other zero at $\delta/\gamma\approx 0.88$.
+No zero transmission exists for $\delta/\gamma\gtrsim 1.08$.
 The anisotropy between $T_{x}$ and $T_{y}$ provides a novel route to engineer subradiant modes in a polarization-selective manner.
-By choosing appropriate detuning $\delta$ and lattice constant $a$, one polarization state can enter a subradiant mode with near-complete reflection, while the orthogonal polarization remains transmissive.
+By choosing appropriate $\delta$ and $a$, one polarization state can enter a subradiant mode with near-complete reflection, while the orthogonal polarization remains transmissive.
 To further illustrate this effect, Fig.
@@
 This demonstrates dual-species atom arrays acting as subwavelength optical polarizers, enabling programmable and independent polarization control at the atomic scale.
-Figure 4: (a–c) Total light intensity and polarization distributions at propagation distances $z=14\lambda$, $50\lambda$, and $90\lambda$, respectively.
+Figure 4: (a–c) Total light intensity and polarization distributions at propagation distances $z/\lambda=14$, $50$, and $90$, respectively.
 The color bar denotes the normalized optical intensity $I=(|E_{x}|^{2}+|E_{y}|^{2})/|E_{0}|^{2}$, where $E_{0}$ is the peak amplitude of the incident Gaussian beam.
@@
 Calculations assume $a/\lambda=0.4$ and $\delta/\gamma=0.5$.
-The figures span spatial ranges $x,y\in\left[-10\lambda,10\lambda\right]$.
-4 Polarization-selective quantuam light modulator
+The figures span spatial ranges $x/\lambda,y/\lambda\in\left[-10,10\right]$.
+4 Polarization-selective quantum light modulator
 By introducing an ancillary atom which can be excited to a Rydberg state, the atom array can be switched between fully reflective and transparent regimes via electromagnetically induced transparency and Rydberg blockade [7, 27].
@@
 Figure 5: (a) Patterned superarray with $2\times 2$ pixels, where each pixel is a dual-species stripe-pattern atom array.
-(b) Total light intensity and polarization distribution at $z=14\lambda$.
+(b) Total light intensity and polarization distribution at $z/\lambda=14$.
 Calculations assume an $N=71\times 71$ atom array with $a/\lambda=0.4$, and a diagonally polarized incident beam with waist $w_{0}=0.3\sqrt{N}a$ and $\delta/\gamma=0.5$.
@@
 Yb atoms feature a transition from $6s6p\penalty 10000\ {}^{3}P_{0}$ to $5d6s\penalty 10000\ {}^{3}D_{1}$ at a telecom-band wavelength $\lambda\approx 1.4\mu\mathrm{m}$, realizing a $J=0\rightarrow J=1$ two-level system, enabling confinement in optical lattices or tweezer arrays with subwavelength spacing.
-We consider a dual-species atom array of the isotopes 171Yb and 173Yb.
+We consider a dual-species array of the isotopes 171Yb and 173Yb.
 Their transition frequencies are $f_{171}=215869124.9\mathrm{MHz}$ and $f_{173}=215869151.9\mathrm{MHz}$, respectively [29].
-Using the spontaneous emission formula $\gamma={d^{2}\omega^{3}}/(3\pi\varepsilon_{0}\hbar c^{3})$, we obtain $\gamma_{171}\approx\gamma_{173}\approx 5.93\mathrm{MHz}$, yielding a separation of $|f_{173}-f_{171}|\sim 4.55\gamma$.
-This energy offset can be continuously tuned via AC Stark shifts induced by coupling to nearby auxiliary levels, allowing controllable detunings $\delta_{A}$ and $\delta_{B}$ within the parameter range considered in this work.
+Using the spontaneous emission formula $\gamma={d^{2}\omega^{3}}/(3\pi\varepsilon_{0}\hbar c^{3})$, we obtain $|f_{173}-f_{171}|\approx 27\mathrm{MHz}\approx 4.55\gamma$.
+This energy offset can be continuously tuned via transition frequency shifts induced by AC Stark effects by trapping the two species at different trap depths (see Supplement 1 for a detailed proposal), enabling controllable detunings within the range $0<\delta<1.2\gamma$ considered in Fig.
+3.
+Moreover, additional calculations show that positional uncertainty of trapped atoms does not compromise experimental observability (see Supplement 1).
 6 Summary
 Our work systematically investigates the cooperative optical response and polarization control mechanisms in dual-species atom arrays.
-By independently tuning the light detunings and lattice spacing, we realize polarization-dependent subradiant modes and identify strong collective scattering at a symmetric detuning point with balanced radiative losses.
+By independently tuning the light detunings and lattice constants, we realize polarization-dependent subradiant modes and identify strong collective scattering at a symmetric detuning point with balanced radiative losses.
 Exploiting this mechanism, the array operates as a subwavelength polarizer with atomic-scale polarization selectivity.
@@
 These results establish dual-species atom arrays as a highly tunable quantum photonic platform and provide a theoretical foundation for advanced quantum light–matter interfaces.
+7 Back matter
 […]